A triangle has sides a=5 and b=4 and angle C=60degrees. Find the length of side c.
1
Expert's answer
2012-01-17T12:45:20-0500
we can use cosine theorem
c^2=a^2+b^2-2*a*b*cos(60)=25+16-2*5*4*1/2=21
so
c=sqrt(21)